# Tree-Based versus Hybrid Graphical-Textual Model Editors: An Empirical Study of Testing Specifications

### v1

@@ 4.3. Simulation-Based Testing
The SESAME tools for simulation-based testing (Partners, 2022, 2023b) provide a model-driven code generation engine. Input models conforming to the DSL configure an evolutionary optimisation loop for a given TestCampaign. The loop optimises a population of fuzzing tests with the intent of maximising safety requirement violations. For every Test generated and added to the model, the SESAME code generation engine generates a test runner, including an MRS logical interface enabling two-way runtime communication. The test runner launches the simulation, monitors and manipulates fuzzing variables to perform the operations defined for that test. When a fuzzing operation is active, all messages upon the target variable will be modified, delayed or deleted as specified by the fuzzing operation. Performance metrics are computed from monitored variables, and their final values are logged under Test as a MetricInstance.
@@
None / Low
6
7
8
17
Medium
6
8
4
3
High
10
7
10
2
@@ 5.1. Recruitment
As the study was part of a wider research project with a broad range of industry and academic partners in modelling platforms, we emailed an invitation letter to target individuals with relevant experience, with a request to indicate their availability for the 2-hour-long study. We shared the same email invite within the Computer Science department of the University of York, for undergraduate and postgraduate students and academics with experience in computer programming, modelling or MDE. Furthermore, the invitation has been shared with various research groups in academia and professionals working in the industry, with modelling expertise.
@@ 5.2. Participants
A total of 22 participants have participated in the study. The participants currently work either in the industry or academia: 5 participants work in the industry (3 companies across the UK, Germany and Romania), and 17 participants come from academia across 8 universities in the UK, Spain, Austria and Germany. Two of the academic participants were computer science students. Participants had a range of years of programming experience, with 10 participants with more than 10 years of experience, 5 participants with 7-10 years of experience, 5 participants with 4-6 years of experience, and 2 participants with less than 1 year. Table 1 provides a detailed view of the level of knowledge and expertise of participants.
@@ 5.3. Procedures
The study was pilot-tested with two users (who did not participate in the final study). During the pilot, we identified practical issues and revised the study accordingly, e.g., to reduce the time needed for the study and improve the wording of task descriptions. Once a participant agreed to join the study, the first author would contact them to arrange a time for the user study. A facilitator, i.e., the first author, guided and observed participants during the sessions. The faculty research committee ethically approved the study.
Each participant joined a virtual meeting with video conferencing together with the facilitator via Zoom. At the beginning of the session, the facilitator described the context of the user study and shared two documents with the participant: an information sheet containing a description of the study and a consent form. The participant read these documents, and then signed the consent form if they agreed to continue with the user study session. Next, the facilitator shared a 20-minute video tutorial, presenting the abstract and concrete syntaxes of the fuzzing-based testing DSL and practical examples of its usage in the Eclipse Integrated Development Environment (IDE) via the hybrid and tree editors. After watching the video, the participants had the opportunity to ask any questions regarding the presented information.
Each participant completed a Google Form survey, which contained the tasks and follow-up questions (e.g. regarding the correctness and confidence in their answers/solutions). Throughout the user study, the facilitator clarified any uncertainties related to the task descriptions. After completing all tasks, participants were asked to report their preferences, including features they liked and disliked about the editors. Lastly, the participants completed a demographic survey, including details of their years of experience and level of expertise in areas such as computer programming, modelling, MDE, DSLs and robotics software.
The participant connected remotely to the computer of the facilitator via TeamViewer. The computer was running a Windows Operating System (OS). On the computer, two Eclipse IDE instances were running for defining and editing testing specifications: one containing the tree editor and another containing the hybrid editor. Participants executed given tasks through both the tree and hybrid editors using the same semantic model. To reduce memorising a task solution, participants were not told that the underlying semantic model is the same.
@@
U1
In the first test, find the Result Value of the Metric Result associated with FuzzingOperationTimesMetric.
U2
In the first test, in the RandomValueFromSetOperation, describe the activation starting condition.
U3
In the campaign result sets of the TestCampaign, find the differences among the last 3 NONDOM sets.
U4
In the testing space, find the variable(s) related to the StreamMetric named collisionOccurrence.
U5
In the testing space, find the fuzzing operation with the largest DoubleRange interval.
M1
In the testing space, create a new TestCampaign and set its name as NewTestCampaign.
M2
Copy the first test from the prior existing TestCampaign into the NewTestCampaign.
M3
In the RandomValueFromSetOperation, modify the activation starting condition by replacing the operators.
M4
In the RandomValueFromSetOperation, change the first referenced variable of the activation starting condition.
M5
Define from scratch the activation ending condition for the RandomValueFromSetOperation.
@@ 5.4. Tasks
Participants were given the tasks from Table 2, categorised into two types. Tasks U1-U5 focus on understanding a testing space, used to evaluate which type of model editor is better for comprehension. Correspondingly, tasks M1-M5 focus on modelling a testing space, used for evaluating which model editor provides better performance when modelling. The table shows a summary of the tasks; the complete task descriptions used in the user study are available in (Predoaia, 2024). All given tasks were driven by the industrial case study, aiming to reflect realistic modelling scenarios performed in practice by domain experts while covering a wide range of modelling activities such as finding specific elements, finding relationships and differences between elements, changing references, and creating and copying elements. Participants operate over the semantic model presented in Figure 4, which contains a subset of the relevant model elements from the actual model used in the case study.
Understanding Tasks (U1 - U5). To perform the tasks using the tree editor, participants have to carefully inspect all model elements, and accordingly expand them to view their child elements and scan their attributes displayed in the Properties View. We describe next how a user should execute the tasks via the hybrid editor.
@@
Task U2 asks participants to describe conditions represented through a tree representation and a textual representation. The task entails viewing the details of the first test as presented above, locating the RandomValueFromSetOperation, finding the activation starting condition, and describing it. In the tree editor, the activation starting condition looks as in Figure 5, whereas in the hybrid editor, it looks as the textual expression displayed in the embedded textual editor from Figure 8. When using the tree editor, the participant was asked to construct an equivalent textual expression as pseudocode. Correspondingly, when using the hybrid editor, the participant was asked to find the textual expression, describe its meaning, and categorise conditions as either BasicCondition or CompositeCondition, and identify the left and right branches of CompositeConditions.
@@
Modelling Tasks (M1 - M5). In the tasks for modelling, participants edit existing model elements, create new ones, and change relevant values of their properties, via both editors. Note that the tasks for modelling are carried out after having completed the tasks for understanding; therefore, the participants have a considerable understanding of the testing space before engaging in these tasks.
@@ 5.5. Data Collection and Analysis
Duration. The duration quantifies the amount of effort required to do a task, measuring the time-on-task, in seconds, within the editors. The duration does not include time spent reading task descriptions, or time wasted when the remote access and control session via TeamViewer was lost. Participants alternated between the Google Form survey and the Eclipse IDE instances containing the editors, thus we were able to monitor the time spent on each.
Correctness. After performing all user study sessions, we examined if participants completed the tasks correctly or made minor/major mistakes. Thus, we marked solutions as either 1 (correct and complete) and 0 (minor mistakes). When a task was not solved correctly, we recorded the specific mistake(s) made.
Errors Corrected. We recorded the mistakes made during modelling, which were spotted and corrected later by users. A mistake represents a performed action that is incorrect, redundant or certainly unrelated to the solution of a modelling task. This metric was collected for task M5, where conditions are modelled from scratch.
@@
Confidence. After performing each task, participants self-reported their confidence level on their solution, ranging from 1 (very low confidence) to 5 (very high confidence).
@@ 6. Results and Discussion
We discuss the results of the user study, considering performance, confidence and preference. Also, we provide aggregated insights and preferences of participants on the usage of the model editors.
@@ 6.1. Performance (RQ1)
Duration. Table 3 shows the absolute differences between the durations for completing the tasks via the tree ($T$) and hybrid ($H$) editors. The table is split into tasks for understanding (U) and modelling (M), and the five numeric columns indicate tasks U1-U5 and M1-M5. The delta ($\Delta$) subtracts the duration when using the hybrid editor from the duration taken with the tree editor, i.e., $Delta=Duration(Tree)-Duration(Hybrid)$. For each task, the delta of the mean and median durations are displayed. A positive value signifies a higher duration with the tree editor compared to the hybrid editor; the opposite holds for a negative value. In addition, Figure 9 shows boxplots with the duration needed for executing the understanding and modelling tasks using both editors.
@@
Type
$\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$
#1 #2 #3 #4 #5
U
Mean
1.55
315.95
-15.32
10
36.45
Median
1.50
301
-23
0.50
32.50
M
Mean
0.59
-9.50
27.09
8.94
9.51
Median
-0.50
2
34
3.50
19.50
Completing task U1 required on average (median) 1.55 (1.50) seconds more with the tree editor than with the hybrid editor. Thus, for solving task U1, which requires locating model elements with specific values of interest, the hybrid editor had slightly better performance; however, no significant difference between the two exists. For task U2, the hybrid editor was faster on average by 315.95 seconds, indicating that using textual representations for expressing conditions improves comprehension significantly. In task U3, the tree editor was faster by 15.32 seconds than the hybrid editor; the median value outlines a higher difference of 23 seconds. Concretely, the relative difference reported to the mean and median durations of the hybrid editor was 15% and 27%, respectively. This shows that tree editors are more appropriate to explore and analyse ordered lists of model elements, as navigating long ordered element lists in diagrams often incurs overheads. For task U4, the hybrid editor was faster by 10 seconds on average, however, this difference is negated by the delta median value, showing the existence of some outliers for this task. We conclude that no significant difference exists in performance between the editors when executing a task that requires finding references between elements. The hybrid editor yielded faster results in task U5, showing that finding elements that meet certain criteria is quicker in diagrams, as users have to meticulously expand and explore all elements in a tree editor.
Considering the modelling tasks, as indicated by the mean and median delta values of tasks M1 and M2, we cannot conclude that there is a difference in performance between the editors for these tasks. For task M3, the hybrid editor resulted in improved performance, indicating that modifying existing conditions via a textual editor with syntax-aware editing features is quicker than through a tree editor. In task M4, the hybrid editor was faster on average by 8.94 seconds, however the delta values of the median suggest a negligible difference between the performance of the editors. Thus, for solving a task that requires changing a reference, the editors have similar performance, although the hybrid editor may be slightly faster. For task M5, the hybrid editor was faster on average (median) by 9.51 (19.50) seconds, with a relative difference of 8% (16%), thus the hybrid editor was more effective in modelling conditions.
@@
Correctness. All participants solved task U1 correctly through the hybrid editor, whereas three participants solved the task incorrectly. From those, two participants reported an incorrect value as a solution, whereas one participant did not find a solution and abandoned the task. For task U2, two participants reported incorrect solutions with the hybrid editor, not being able to accurately state the left and right branches of the conditions represented textually. Correspondingly, seven participants solved the task incorrectly with the tree editor, as they constructed incorrect textual representations of the conditions by either grouping sub-conditions incorrectly, missing operators or using unsuitable operators. All participants solved correctly tasks U3-U5 using both editors. The results evidence that the hybrid editor surpasses the tree editor in correctly locating specific model elements and in comprehending conditions.
In modelling tasks, one participant incorrectly performed task M1 via the hybrid syntax. The participant tried to set the name of a TestCampaign by double-clicking its name on the diagram, and then entering the name in the pop-up menu displayed after double-clicking. The participant forgot a detail shown in the video tutorial: double-clicking specific elements in the hybrid editor instantiates new diagrams. As such, the participant instantiated a new diagram with the name entered in the pop-up menu, without noticing that the name of the TestCampaign had not been set. Correspondingly, all participants solved task M1 correctly through the tree editor.
@@
$\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$
Clicks M3 Keystrokes M3 Clicks M5 Keystrokes M5
Mean
24.32
-32.55
29.70
-53.49
Median
25
-14.50
30
-48
@@
Type
$\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$
#1 #2 #3 #4 #5
U
Mean
-0.68
-1.18
-0.14
0.03
-0.05
Median
-0.50
-1
0
0
0
M
Mean
-0.09
0.09
-0.18
-0.34
-0.32
Median
0
0
0
0
0
@@ 6.2. Confidence (RQ2)
We answer research question RQ2 concerning the users’ confidence when using both editors to do the specified tasks by analysing their self-reported confidence level, ranging between 1 (very low confidence) to 5 (very high confidence). Table 5 shows statistics of the confidence level of participants about their task solutions. As before, the table is split into tasks for understanding (U) and modelling (M), and the five numeric columns are linked to tasks U1-U5 and M1-M5. The delta subtracts the mean and median confidence when using the hybrid editor from the mean and median confidence when using the tree editor. For task U1, participants were on average slightly more confident with their solution when using the hybrid editor, signifying users’ confidence for tasks that require understanding a domain and locating various model elements with specific values of interest. Participants were on average more confident by 1.18 points, with their solution to task U2 when using the hybrid editor, signifying their confidence in understanding the meaning of conditions. For the remaining tasks, the results have no significant differences in the confidence of the solutions.
We also analyse the correlation between incorrect solutions and confidence level, aiming to determine how confident participants are with their incorrect solutions. A mean confidence of 3.13 and a median confidence of 3 has been reported for the incorrect solutions to tasks U1-U5 when using the tree editor. This shows that participants are neutral about the correctness of their partially incorrect solutions. Further, a mean and median confidence of 4.5 has been reported for the incorrect solutions to tasks M1-M5 when using the tree editor. Similarly, a mean confidence of 4.67 and a median confidence of 5 has been reported for the incorrect solutions to tasks M1-M5 when using the hybrid editor. These results show that participants were confident with their partially incorrect solutions, as they had not noticed any faults. This finding suggests that the confidence level is related to the ability of users to spot their mistakes or inadequacies. Participants realised that they could not solve tasks that verify their understanding, but they did not observe some of their mistakes in tasks that required modelling.

### v2

@@ 4.3. Simulation-Based Testing
The SESAME tools for simulation-based testing (Partners, 2022, 2023b) provide a model-driven code generation engine. Input models conforming to the DSL configure an evolutionary optimisation loop for a given TestCampaign. The loop optimises a population of fuzzing tests with the intent of maximising safety requirement violations. For every Test generated and added to the model, the SESAME code generation engine generates a test runner, including an MRS logical interface enabling two-way runtime communication. The test runner launches the simulation, monitors and manipulates fuzzing variables to perform the operations defined for that test. Performance metrics are computed from monitored variables, and their final values are logged under Test as a MetricInstance.
@@
None / Low 6 7 8 17
Medium 6 8 4 3
High 10 7 10 2
@@ 5.1. Recruitment
As the study was part of a wider research project with a broad range of industry and academic partners in modelling platforms, we emailed an invitation letter to target individuals with relevant experience, with a request to indicate their availability for the 2-hour-long study. We shared the same email invite within the Computer Science department of the University of York, UK, for undergraduate and postgraduate students and academics with experience in computer programming, modelling or MDE. The invitation was also shared with various research groups in academia and professionals working in the industry, with modelling expertise.
@@ 5.2. Participants
A total of 22 participants participated in the study. The participants currently work either in the industry or academia: 5 participants work in the industry (3 companies across the UK, Germany and Romania), and 17 participants come from academia across 8 universities in the UK, Spain, Austria and Germany. Two of the academic participants were computer science students. Participants had a range of years of programming experience, with 10 participants with more than 10 years, 5 participants with 7–10 years, 5 participants with 4–6 years, and 2 participants with less than 1. Table 1 shows the level of knowledge and expertise of participants.
@@ 5.3. Procedures
The study was pilot-tested with two users who did not participate in the final study. During the pilot, we identified practical issues and revised the study accordingly, e.g., to reduce the time needed for the study and improve the wording of task descriptions. Once a participant agreed to join the study, the first author would contact them to arrange a time for the user study. A facilitator, i.e., the first author, guided and observed participants during the sessions. The faculty research committee ethically approved the study.
Each participant joined a virtual meeting with video conferencing together with the facilitator via Zoom. At the beginning of the session, the facilitator described the context of the user study and shared two documents with the participant: an information sheet describing the study and a consent form. The participant read these documents and then signed the consent form if they agreed to continue with the user study session. Next, the facilitator shared a 20-minute video tutorial, presenting the abstract and concrete syntaxes of the fuzzing-based testing DSL and practical examples of its usage in the Eclipse Integrated Development Environment (IDE) via the hybrid and tree editors. After watching the video, the participants could ask questions on the presented information.
Each participant completed a Google Forms survey, which contained the tasks and follow-up questions (e.g., regarding the correctness and confidence in their answers/solutions). Throughout the user study, the facilitator clarified any uncertainties related to the task descriptions. After completing all tasks, participants were asked to report their preferences, including features they liked and disliked about the editors. Lastly, the participants completed a demographic survey, including details of their years of experience and level of expertise in the areas of Table 1.
The participant connected remotely to the computer of the facilitator via TeamViewer. The computer ran a Windows Operating System (OS). On the computer, two Eclipse IDE instances were running for defining and editing testing specifications: one containing the tree editor and another containing the hybrid editor. Participants executed given tasks through both editors using the same semantic model. To reduce memorising a task solution, participants were not told that the underlying semantic model is the same.
@@
U1 In the first test, find the Result Value of the Metric Result associated with FuzzingOperationTimesMetric.
U2 In the first test, in the RandomValueFromSetOperation, describe the activation starting condition.
U3 In the campaign result sets of the TestCampaign, find the differences among the last 3 NONDOM sets.
U4 In the testing space, find the variable(s) related to the StreamMetric named collisionOccurrence.
U5 In the testing space, find the fuzzing operation with the largest DoubleRange interval.
M1 In the testing space, create a new TestCampaign and set its name as NewTestCampaign.
M2 Copy the first test from the prior existing TestCampaign into the NewTestCampaign.
M3 In the RandomValueFromSetOperation, modify the activation starting condition by replacing the operators.
M4 In the RandomValueFromSetOperation, change the first referenced variable of the activation starting condition.
M5 Define from scratch the activation ending condition for the RandomValueFromSetOperation.
@@ 5.4. Tasks
Participants were given the tasks from Table 2, categorised into two types. Tasks U1U5 focus on understanding a testing space, used to evaluate which type of model editor is better for comprehension. Tasks M1M5 focus on modelling a testing space, used for evaluating which model editor provides better performance when modelling. The table shows a summary of the tasks; the complete task descriptions used in the study are available in (Predoaia, 2024). All given tasks were driven by the industrial case study, aiming to reflect realistic modelling scenarios performed in practice by domain experts while covering a wide range of modelling activities such as finding specific elements, finding relationships and differences between elements, changing references, and creating and copying elements. Participants operate over the semantic model presented in Figure 4, which contains a subset of the relevant model elements from the actual model used in the case study. All tasks of the same type were first executed with one editor, and then they were executed again with the other editor, e.g., after completing all modelling tasks with the hybrid editor, they were executed again with the tree editor.
Understanding Tasks (U1 – U5). To perform the tasks using the tree editor, participants have to carefully inspect all model elements, and accordingly expand them to view their child elements and scan their attributes displayed in the Properties View. We describe next how a user should execute the tasks via the hybrid editor.
@@
Task U2 asks participants to describe conditions defined through a tree representation and a textual representation. The task entails viewing the details of the first test as presented above, locating the RandomValueFromSetOperation, finding the activation starting condition, and describing it. In the tree editor, the activation starting condition looks as in Figure 5, whereas in the hybrid editor, it looks as the textual expression displayed in the embedded textual editor from Figure 8. When using the tree editor, the participant was asked to construct an equivalent textual expression as pseudocode. Correspondingly, when using the hybrid editor, the participant was asked to find the textual expression, describe its meaning, and categorise conditions as either BasicCondition or CompositeCondition, and identify the left and right branches of CompositeConditions.
@@
Modelling Tasks (M1 – M5). In these tasks, participants edit existing and create new model elements, and change relevant property values via both editors. The modelling tasks are done after completing the understanding tasks; thus, participants have a good understanding of the testing space before engaging in these tasks.
@@ 5.5. Data Collection and Analysis
Duration. The duration quantifies the amount of effort required to do a task, measuring the time-on-task, in seconds, within the editors. The duration does not include time spent reading task descriptions, or time wasted when the remote access and control session via TeamViewer was lost. Participants alternated between the Google Forms survey and the Eclipse IDE instances containing the editors, thus we were able to monitor the time spent on each.
Correctness. After performing all user study sessions, we examined whether participants completed the tasks correctly or made mistakes. Thus, we marked solutions as either 1 (correct and complete) or 0 (mistakes). When a task was not solved correctly, we recorded the specific mistake(s) made.
@@
Confidence. After performing each task, participants self-reported their confidence level in their solution, ranging from 1 (very low confidence) to 5 (very high confidence).
Errors Corrected. We recorded the mistakes made in the context of a modelling task, which were later corrected by the user during the execution of the same task and before its conclusion. A mistake represents a performed action that is incorrect, redundant or certainly unrelated to the solution of a modelling task. Note that the clicks and keystrokes and the time spent while making mistakes and correcting them were included in the analysis.
@@ 6.1. Performance (RQ1)
Duration. Table 3 shows the absolute differences between the durations for completing the tasks via the tree ($T$) and hybrid ($H$) editors. The table is split into tasks for understanding (U) and modelling (M), and the five numeric columns indicate tasks U1U5 and M1M5. The delta ($\Delta$) subtracts the duration when using the hybrid editor from the duration taken with the tree editor, i.e., $Delta=Duration(Tree)-Duration(Hybrid)$. For each task, the delta of the mean and median durations are displayed. A positive value signifies a higher duration with the tree editor compared to the hybrid editor; the opposite holds for a negative value. In addition, Figure 9 shows boxplots with the duration needed for executing the understanding and modelling tasks using both editors.
@@
Type $\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$1 #2 #3 #4 #5
U Mean 1.55 315.95 -15.32 10 36.45
Median 1.50 301 -23 0.50 32.50
M Mean 0.59 -9.50 27.09 8.94 9.51
Median -0.50 2 34 3.50 19.50
Completing task U1 required on average (median) 1.55 (1.50) seconds more with the tree editor than with the hybrid editor. Thus, for solving task U1, which requires locating model elements with specific values of interest, the hybrid editor had slightly better performance; however, no significant difference between the two exists. For task U2, the hybrid editor was faster on average by 315.95 seconds, indicating that using textual representations for expressing conditions improves comprehension significantly. In task U3, the tree editor was faster by 15.32 seconds than the hybrid editor; the median value outlines a higher difference, of 23 seconds. Concretely, the relative difference reported to the mean and median durations of the hybrid editor was 15% and 27%, respectively. This shows that tree editors are more appropriate to explore and analyse ordered lists of model elements, as navigating long ordered element lists in diagrams often incurs overheads. For task U4, the hybrid editor was faster by 10 seconds on average, however, this difference is negated by the delta median value, showing the existence of some outliers for this task. We conclude that no significant difference exists in performance between the editors when executing a task that requires finding references between elements. The hybrid editor yielded faster results in task U5, showing that finding elements that meet certain criteria is quicker in diagrams, as users have to meticulously expand and explore all elements in a tree editor.
Considering the modelling tasks, as indicated by the mean and median delta values of tasks M1 and M2, we cannot conclude that there is a difference in performance between the editors for these tasks. For task M3, the hybrid editor resulted in improved performance, indicating that modifying existing conditions via a textual editor with syntax-aware editing features is quicker than through a tree editor. In task M4, the hybrid editor was faster on average by 8.94 seconds, however, the delta value of the median suggests a negligible difference between the performance of the editors. Thus, for solving a task that requires changing a reference, the editors have similar performance, although the hybrid editor may be slightly faster. For task M5, the hybrid editor was faster on average (median) by 9.51 (19.50) seconds, with a relative difference of 8% (16%), thus the hybrid editor was more effective in modelling conditions.
@@
Correctness. All participants solved task U1 correctly through the hybrid editor, whereas three participants solved the task incorrectly with the tree editor. From those, two participants reported an incorrect value as a solution, whereas one participant did not find a solution and abandoned the task. For task U2, two participants reported incorrect solutions with the hybrid editor, not being able to accurately state the left and right branches of the conditions represented textually. Correspondingly, seven participants solved the task incorrectly with the tree editor, as they constructed incorrect equivalent textual representations of the conditions from the tree editor by either grouping sub-conditions incorrectly, missing operators or using unsuitable operators. All participants solved correctly tasks U3U5 using both editors. The results evidence that the hybrid editor surpasses the tree editor in correctly locating specific model elements and in comprehending conditions.
In modelling tasks, one participant incorrectly performed task M1 via the hybrid editor. The participant tried to set the name of a TestCampaign by double-clicking its name on the diagram, and then entering the name in the pop-up menu displayed after double-clicking. The participant forgot a detail shown in the video tutorial: double-clicking specific elements in the hybrid editor instantiates new diagrams. As such, the participant instantiated a new diagram with the name entered in the pop-up menu, without noticing that the name of the TestCampaign had not been set. Correspondingly, all participants solved task M1 correctly through the tree editor.
@@
$\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$ Clicks M3 Keystrokes M3 Clicks M5 Keystrokes M5
Mean 24.32 -32.55 29.70 -53.49
Median 25 -14.50 30 -48
@@
Type $\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$1 #2 #3 #4 #5
U Mean -0.68 -1.18 -0.14 0.03 -0.05
Median -0.50 -1 0 0 0
M Mean -0.09 0.09 -0.18 -0.34 -0.32
Median 0 0 0 0 0
@@ 6.2. Confidence (RQ2)
We answer research question RQ2 concerning the users’ confidence when using both editors to do the specified tasks by analysing their self-reported confidence level, ranging between 1 (very low confidence) to 5 (very high confidence). Table 5 shows statistics of the confidence level of participants about their task solutions. As before, the table is split into tasks for understanding (U) and modelling (M), and the five numeric columns are linked to tasks U1U5 and M1M5. The delta subtracts the mean and median confidence when using the hybrid editor from the mean and median confidence when using the tree editor. For task U1, participants were on average slightly more confident with their solution when using the hybrid editor, signifying users’ confidence for tasks that require understanding a domain and locating various model elements with specific values of interest. Participants were on average more confident by 1.18 points, with their solution to task U2 when using the hybrid editor, signifying their confidence in understanding the meaning of conditions. For the remaining tasks, the results show no significant differences in the confidence of the solutions.
We also analyse the correlation between incorrect solutions and confidence level, aiming to determine how confident participants are with their incorrect solutions. A mean confidence of 3.13 and median confidence of 3 has been reported for the incorrect solutions to tasks U1U5 when using the tree editor. This shows that participants are neutral about the correctness of their partially incorrect solutions. Further, a mean and median confidence of 4.5 has been reported for the incorrect solutions to tasks M1M5 when using the tree editor. Similarly, a mean confidence of 4.67 and a median confidence of 5 has been reported for the incorrect solutions to tasks M1M5 when using the hybrid editor. These results show that participants were confident with their partially incorrect solutions, as they had not noticed any of their mistakes. This finding suggests that the confidence level is related to the ability of users to spot their mistakes. Participants could identify the understanding tasks they could not solve, and that was reflected in their confidence level for each understanding task. However, for the modelling tasks, the participants wrongly assumed that their solutions were correct, leading to a mistaken increased confidence level.